OpenStudy (anonymous):

The diagonals of rhombus WXYZ intersect at point P. If WP = 12 and XP = 5, find the perimeter of the rhombus.

OpenStudy (anonymous):

each side =sqrt(12^2+5^2)=sqrt(144+25)=sqrt(169)=13 hence perimeter= 4*13=52